One card is selected at random from a deck of cards. Determine the probability that the card selected is a diamond.
Found 2 solutions by jim_thompson5910, oberobic:
Answer by jim_thompson5910(35256)

P(Diamond) = (# of diamonds)/(# of cards total)


P(Diamond) = 13/52


P(Diamond) = 1/4

Answer by oberobic(2304)

Assume there are 52 cards in the deck.
13 of these are diamonds.
.
There are 13 favorable outcomes.
There are 52 possible outcomes.
.
13/52 = 1/4 = 25%
.
Done.